# exa-AMD: An Exascale-Ready Framework for Accelerating the Discovery and Design of Functional Materials

### v1

@@ Abstract
Exascale computing is transforming the field of materials science by enabling simulations of unprecedented scale and complexity. We present exa-AMD, an open-source, high-performance simulation code specifically designed for accelerated materials discovery on modern supercomputers. exa-AMD addresses the computational challenges inherent in large-scale materials discovery by employing task-based parallelization strategies and optimized data management tailored for high performance computers. The code features a modular design, supports both distributed and on-node parallelism, and is designed for flexibility and extensibility to accommodate a wide range of materials science applications. We detail the underlying algorithms and implementation, and provide comprehensive benchmark results demonstrating strong scaling across multiple high performance computing platforms. We provide two example applications, the design of Fe-Co-Zr and Na-B-C compounds, to illustrate the code’s effectiveness in accelerating the discovery and characterization of novel materials. With only a set of elements as input, exa-AMD automates the workflow on CPU or GPU-enabled clusters, outputs the structures and energies of promising candidates, and updates the phase diagram. exa-AMD is publicly available on GitHub, with detailed documentation and reproducible test cases to support community engagement and collaborative research. This work aims to advance materials science by providing a robust, efficient, and extensible tool ready for exascale platforms.
@@ 1 Introduction
Traditional high-throughput methods, primarily based on density functional theory (DFT), are used to establish large materials databases such as the Materials Project [1], Automatic FLOW for Materials Discovery (AFLOW) [2], and the Open Quantum Materials Database (OQMD) [3, 4]. These approaches focus on computing key properties such as formation energies, band gaps, and elastic moduli for known or enumerated structures, providing valuable insight into phase stability and guiding experimental efforts [1, 2, 3, 4, 5, 6]. However, they are severely limited in coverage and efficiency for complex systems, as brute-force DFT evaluations become computationally intractable for millions of candidates, often requiring months of supercomputing time and lacking the flexibility to intelligently prioritize low-energy structures without exhaustive calculations.
@@
A solution to this problem is the use of machine learning (ML) models to rapidly predict formation energies and stability metrics for novel materials and down-select candidates before performing costly DFT validation, thereby accelerating the exploration of significantly larger chemical and structural spaces. ML models such as graph neural networks have shown promises in achieving DFT accuracy at a fraction of the cost, reducing screening time from months to minutes while maintaining predictive reliability [7, 8]. This data-driven acceleration is essential for tackling scientific problems like designing rare-earth-free magnets or novel battery materials, where multi-objective optimization across stability, magnetism, and synthesizability is required.
@@
In this work, we present exa-AMD, an open-source, modular, and scalable framework tailored for Accelerated Materials Discovery on exascale platforms. exa-AMD unifies data mining, machine learning, advanced workflow automation, and first-principles computation into a holistic framework, enabling rapid prediction of stable and metastable compounds and property optimization in complex, multi-element chemical spaces that are otherwise inaccessible through conventional empirical or brute-force computational methods. It supports ternary and quaternary alloys, and can be extended to higher-order systems. exa-AMD integrates advanced ML models and high-throughput quantum mechanical calculations to enable rapid screening and characterization of novel compounds through dynamical computing resource management. The software streamlines the entire materials discovery process, from structural hypothesis generation to thermodynamic stability and property prediction, thereby significantly reducing the time and resources needed for discovering promising candidates. The code’s flexible, modular design supports various workflow needs for researchers to efficiently target new compounds for synthesis and technological applications, as will be demonstrated by two examples in designing rare-earth-free magnets and beyond. We also show benchmarking results to demonstrate its excellent scaling on some of the largest high performance computing (HPC) resources available.
@@ 2 Program Overview
The exa-AMD framework is designed as a modular, high-throughput pipeline for accelerated materials discovery, automating the entire process from user-specified chemical elements to the prediction of stable compounds and phase diagrams. Implemented in Python, it leverages ML for rapid stability screening and DFT for precise characterization, while ensuring extensibility, reproducibility, and efficient scaling from laptops to GPU-rich HPC environments. The basic workflow (Fig. 1) is comprised of five major steps: (1) crystal structures construction; (2) rapid stability screening using ML models; (3) duplicate removal and structure selection; (4) first-principles DFT optimization and characterization of physical properties; (5) post-processing evaluates formation energies relative to elemental references, assesses thermodynamic stability via convex hull construction, and generates updated phase diagrams.
Each of these stages is encapsulated as a configurable module, allowing users to adapt to new chemical systems, swap ML models, or integrate alternative quantum calculation packages, with workflow orchestration handled by Parsl [9] for dynamic resource management. In the following, we will describe each step in details.
Figure 1: Schematic workflow.
@@ 2.1 Crystal structures construction
At the beginning of the framework, an initial pool of crystal structures in the Crystallographic Information File (CIF) format should be provided. They are the “seeds” which the new materials will be generated from. These crystal structure files can be obtained from existing databases (Materials Project [1], GNoME [10], OQMD [3, 4], AFLOW [2], and NovoMag [11], ensuring a broad coverage of known and plausible motifs. By default, exa-AMD provides an initial pool of X number of structures. Users can also prepare their own prototype crystal structures as their customized initial structure pool of their research interest.
Hundreds of thousands to nearly one million candidate structures are then generated using three methods: (i) systematic elemental substitution, where … (ii) combinatorial atom-type shuffling from these prototype structure [12, 13, 14], where … (iii) lattice-volume scaling, which allows for the sampling of realistic bond lengths across different chemistries, typically with a factor of 0.9–1.1 compared to the reference bond lengths, as described in [14, 12, 13].
@@ 2.2 Rapid stability screening using ML models
To efficiently predict the formation energies and eliminate high-energy candidates, exa-AMD employs an ML model, Crystal Graph Convolutional Neural Network (CGCNN), for high-throughput stability screening [7]. CGCNN represents each hypothetical crystal as a graph, encoding atoms as nodes (featuring atomic number and chemical environment) and bonds as edges (including interatomic distances). Each hypothetical structure generated from the previous step is evaluated using the CGCNN model that is trained to predict the formation energy. The initial, universal CGCNN model we utilized in this framework was trained based on XXX structures from Materials Project, as described in  [7]. The mean absolute error of this model is typically around 0.1–0.2 eV/atom. Down selection is then carried out by thresholding the formation energies to include low-energy configurations (e.g., $E_{\mathrm{f}}<0$ eV/atom) and removing duplicates using structure similarity criteria. This results in typically 1,000–4,000 unique candidates for first-principles calculations in the next step [12, 13, 14].
Our previous work show that the use of ML models drastically reduces the computational time from months (using brute force DFT) to minutes, for over $10^{6}$ structures . This screening enables the rapid exclusion of high-energy or chemically implausible candidates before any quantum calculation is performed.
After the completion of the first round of the framework (i.e., including DFT calculations and post-processing as discussed in Sections 2.3 and 2.4), a system-specific, second-generation CGCNN model is retrained using several hundred to thousands of new DFT-relaxed ground states, reducing the mean absolute errors to as low as 0.03–0.05 eV/atom for the chemical system we investigated [15, 16, 12, 13, 17]. Key hyperparameters for training include 3–6 convolutional layers, batch size of 64–256, 100–200 epochs, stochastic gradient decent optimization, and 80/10/10 data splits for training/validation/testing, respectively. Since the evaluation of each candidate structure is independent, this approach provides a scalable and accurate procedure to select the most promising candidate materials for the costly quantum calculations in the next step. Importantly, our framework allows users to replace or augment this model with other ML architectures or retrained models specific to their system of interest, enabling improved prediction accuracy and tailored screening strategies. Such flexibility allows adaptation to a wide range of material classes and accelerates convergence to plausible candidate structures.
@@ 2.3 First-Principles Calculations
Formation energies per atom are always referenced to the relaxed energies of the elemental (and, if needed, binary) phases, and convex hull constructions are used to assess thermodynamic stability—a compound is designated as stable, metastable ($E_{\rm hull}<0.1$ eV/atom), or unstable in this step. For correlated systems, a Hubbard $U$ correction is applied within standard DFT for more accurate valence treatments when necessary. Dynamical stability checks, such as phonon calculations and ab initio molecular dynamics, can be conducted for representative new phases as post-processing steps. By parallelizing DFT jobs across CPU/GPU clusters using Parsl, thousands of structure relaxations with varying size and complexity can be completed within hours to days—enabling practical exploration of otherwise intractable structural and compositional spaces [14, 12, 13, 15, 16]. Although this first-principles calculation stage primarily uses VASP as the default DFT engine, the workflow is designed to be agnostic about the choice of DFT software. Depending on user preferences and available computational resources, alternative software packages such as Quantum ESPRESSO [22, 23] can be used. This modularity not only offers flexibility in computation but also enables benchmarking and method comparison to ensure robust validation of candidate materials.
@@ 2.4 Convex Hull and Thermodynamic Stability
$E_{\text{form}}=E_{\text{tot}}(\text{compound})-\sum_{i}n_{i}E_{\text{ref}}(i)$ (1)
@@
Candidates are compared against all known phases to compute their energy above the convex hull ($E_{\mathrm{hull}}$). Compounds on the hull ($E_{\mathrm{hull}}=0$ eV/atom) or near the hull (e.g., $E_{\mathrm{hull}}<0.1$ eV/atom) are considered stable or metastable, respectively. This rigorous convex hull analysis is essential for predicting which compounds can be synthesized in practice, as confirmed by literature and data-mined studies .
It is important to note that the scope of the predicted phases and phase diagrams generated by exa-AMD is fundamentally determined by the set of structure prototypes supplied at the beginning of the workflow. These initial prototypes, either provided by default with exa-AMD or customized by the user, define the structural motifs accessible to elemental substitution, combinatorial screening, and all subsequent machine learning and DFT evaluation. Consequently, all new compounds, low-energy structures, and phase diagram updates produced by the current version of exa-AMD are restricted to the chemical and structural space spanned by these templates. While this approach enables high throughput and chemical flexibility, it does not guarantee exhaustive exploration of all possible structure types. As demonstrated in our previous work on the Fe-Co-C system [17], this limitation can be addressed by integrating adaptive genetic algorithms (AGA) [24, 25] capable of exploring and predicting new structure types that are absent from existing databases or user-supplied templates. In future releases of exa-AMD, we plan to natively incorporate AGA-based structure prediction to overcome the prototype bottleneck and expand the framework’s reach to uncover genuinely novel structural motifs and compounds beyond the current prototype-driven paradigm.
@@ 3 Software Implementation
exa-AMD is implemented in Python and designed for clarity, portability, and extensibility. It encompasses plug-in scripts for structure handling, database access, machine learning inference, and DFT job submission. The use of the Parsl library [9] ensures efficient, scalable workflow orchestration involving task parallelism, fault tolerance, and elastic resource management. Users can configure resource allocation, execution platforms, and job scheduling via flexible configuration files, allowing seamless portability from local workstations to large-scale HPC systems. Key features of the software implementation include:
@@
    Modular design: The framework’s modular architecture allows users to customize or replace key components to adapt to diverse research needs. For example, users can provide their own customized initial structure pool as discussed above. Moreover, the machine learning model (default is CGCNN currently), can be replaced by other advanced models [add references here]. The first principles component also supports multiple DFT engines such as VASP, Quantum ESPRESSO, and others.
@@
    I/O and database integration: Structures are sourced from our exa-AMD dataset [26], curated from multiple databases or user-provided dataset, with outputs systematically organized into logs, results, and configuration files to ensure comprehensive data provenance and reproducibility.
@@
    Parallel execution: High-throughput tasks like ML screening and DFT calculations are automatically parallelized, with configurable resource allocations (node types and counts) for elastic scaling on supercomputers and clouds.
@@
At the heart of exa-AMD’s scalability and automation is the use of Parsl [9], a flexible, Python-based parallel programming and workflow library designed to efficiently orchestrate scientific pipelines across heterogeneous computational resources. Parsl enables exa-AMD to decompose each major workflow step—structure generation, ML-based screening, structure similarity filtering, DFT calculations, and property post-processing—into fine-grained, asynchronous tasks (“apps” in Parsl terminology) that can be dynamically mapped onto available compute resources such as CPUs, GPUs, or hybrid clusters. In practice, each workflow stage is implemented as a Parsl app, either a “Python app” or a “Bash app”. The user specifies execution back-ends (e.g., SLURM, PBS, or local), resource allocation (cores, GPUs, wall time), and queue preferences in a standard Parsl configuration file, which is agnostic about the scientific workflow. Parsl’s executor-provider abstraction allows exa-AMD to submit thousands of jobs to HPC queueing systems and efficiently manage task distribution, resource scaling, and result collection without manual intervention. For example, ML screening tasks are offloaded as batches to available GPUs, while high-throughput DFT tasks are partitioned into separate job arrays, each tracked and managed independently. As tasks complete, outputs are automatically aggregated and passed to downstream stages, ensuring full automation and rapid throughput.
@@
Parsl also enables robust resumability and error handling within exa-AMD. Every task’s state and results are tracked, allowing failed calculations to be retried or workflows to resume from intermediate stages in case of interruption. Through elastic scaling, Parsl dynamically grows or shrinks the compute pool depending on job queue length and available resources; for example, releasing idle GPU nodes during less demanding workflow stages. This highly modular and fault-tolerant approach not only maximizes resource usage and throughput on shared HPC queues but also supports rapid development and reproducibility—key requirements for cutting-edge, community-driven materials discovery frameworks.
@@
By decoupling workflow logic from execution configuration, exa-AMD with Parsl facilitates portability from laptops to exascale clusters. Users can port identical workflows across computing systems by simply adjusting configuration files, making rapid prototyping, exploration, and full production campaigns equally straightforward. Extensive documentation and tutorials are provided on the exa-AMD documentation website, including ready-to-use Parsl configuration templates for common platforms and guidance for workflow customization.

### v2

@@ Abstract
We present exa-AMD, an open-source, high-performance framework designed for accelerated materials discovery on modern supercomputers. exa-AMD overcomes key computational bottlenecks in large-scale structure prediction through task-based parallelization, adaptive load balancing, and optimized data management for CPU and GPU architectures. The framework automates the end-to-end workflow—from generating candidate structures to evaluating formation energies and updating phase diagrams. Its modular design allows users to easily replace or extend components with custom machine learning models, alternative initial structure templates, and future structure generators, enabling flexible integration with emerging AI approaches. We demonstrate strong scaling across high-performance computing platforms and highlight applications to Na–B–C, Ce–Co–B, and Fe–Co–Zr systems, establishing exa-AMD as a robust and exascale-ready tool for accelerating the discovery and design of functional materials. exa-AMD is publicly available on GitHub, with detailed documentation and reproducible test cases to support community engagement and collaborative research.
@@ 1 Introduction
Traditional high-throughput methods, primarily based on density functional theory (DFT), are used to establish large materials databases such as the Materials Project [MP], Automatic FLOW for Materials Discovery (AFLOW) [AFLOW], and the Open Quantum Materials Database (OQMD) [OQMD, OQMD2]. These approaches focus on computing key properties such as formation energies, band gaps, and elastic moduli for known or enumerated structures, providing valuable insight into phase stability and guiding experimental efforts [MP, AFLOW, OQMD, OQMD2, high-throughput-review1, high-throughput-review2]. However, they are severely limited in coverage and efficiency for complex systems, as brute-force DFT evaluations become computationally intractable for millions of candidates, often requiring months of supercomputing time and lacking the flexibility to intelligently prioritize low-energy structures without exhaustive calculations.
@@
A solution to this problem is the use of machine learning (ML) models to rapidly predict formation energies and stability metrics for novel materials and down-select candidates before performing costly DFT validation, thereby accelerating the exploration of significantly larger chemical and structural spaces. ML models such as graph neural networks have shown promises in achieving DFT accuracy at a fraction of the cost, reducing screening time from months to minutes while maintaining predictive reliability [CGCNN, ML-materials-review]. This data-driven acceleration is essential for tackling scientific problems like designing rare-earth-free magnets or novel battery materials, where multi-objective optimization across stability, magnetism, and synthesizability is required.
@@
In this work, we present exa-AMD, an open-source, modular, and scalable framework for Accelerated Materials Discovery targeted to exascale platforms. exa-AMD unifies data mining, machine learning, advanced workflow automation, and first-principles computation into a holistic framework, enabling rapid prediction of stable and metastable compounds and property optimization in complex, multi-element chemical spaces that are otherwise inaccessible through conventional empirical or brute-force computational methods. It supports ternary and quaternary alloys, and can be extended to higher-order systems. exa-AMD integrates advanced ML models and high-throughput quantum mechanical calculations to enable rapid screening and characterization of novel compounds through dynamical computing resource management. The software streamlines the entire materials discovery process, from structural hypothesis generation to thermodynamic stability and property prediction, thereby significantly reducing the time and resources needed for discovering promising candidates. The code’s flexible, modular design supports various workflow needs for researchers to efficiently target new compounds for synthesis and technological applications, as will be demonstrated by two examples in designing rare-earth-free magnets and beyond. We also show benchmarking results to demonstrate its excellent scaling on some of the largest high performance computing (HPC) resources available.
@@ 2 Program Overview
The exa-AMD framework is designed as a modular, high-throughput pipeline for accelerated materials discovery, automating the entire process from user-specified chemical elements to the prediction of stable compounds and phase diagrams. Implemented in Python, it leverages ML for rapid stability screening and DFT for precise characterization, while ensuring extensibility, reproducibility, and efficient scaling from laptops to GPU-rich HPC environments. The basic workflow is comprised of five major steps as shown in Fig. 1: (1) crystal structures construction; (2) formation energy prediction using ML models; (3) duplicate removal and structure selection; (4) first-principles calculations and characterization of physical properties; (5) post-processing which assesses thermodynamic stability via convex hull construction, and generates updated phase diagrams.
Figure 1: Schematic workflow consists of five major steps: (1) crystal structures construction; (2) formation energy prediction using ML models; (3) structure selection; (4) first-principles calculations; (5) post-processing.
Each of these stages is encapsulated as a configurable module, allowing users to adapt to new chemical systems, swap ML models, or integrate alternative quantum calculation packages, with workflow orchestration handled by Parsl [Parsl] for dynamic resource management. In the following, we will describe each step in details.
@@ 2.1 Crystal structures construction
At the beginning of the framework, an initial pool of crystal structures in the Crystallographic Information File (CIF) format should be provided. They are the “seeds” from which the new materials will be generated. These crystal structure files can be obtained from existing databases (Materials Project [MP], GNoME [GNOME], OQMD [OQMD, OQMD2], AFLOW [AFLOW], and NovoMag [Novomag], ensuring a broad coverage of known and plausible motifs. By default, exa-AMD provides an initial pool of 36553 unique ternary structures, and 5254 unique quaternary structures. Users can also prepare their own prototype crystal structures as their customized initial structure pool of their research interest.
Hundreds of thousands to nearly one million candidate structures are then generated using two complementary methods, visually represented in Fig. 2:
(a) Element Substitution.
(b) Lattice-Volume Scaling.
Figure 2: Illustration of the two primary high-throughput structure generation methods. (a) Element substitution: systematic replacement of atomic species at fixed crystallographic sites. (b) Lattice-volume scaling: uniform expansion or contraction of the unit cell volume by applying scaling factors.
(i) Systematic elemental substitution
Systematic elemental substitution with combinatorial atom-type shuffling is used to explore compositional and configurational space (Fig. 2a). In this process, each prototype structure undergoes the systematic replacement of its atomic species with the user-specified target elements, while simultaneously permuting the element assignments across available crystallographic sites [FeCoZr, CeCoCu, JOSS]. For example, for a ternary template structure A-B-C, and the user specifies target elements X, Y, and Z, the code generates all chemically feasible combinations (e.g., X-Y-Z, X-X-Z, Y-Y-Z, etc.) by substituting each site with each target element. This approach ensures comprehensive coverage of both compositional space and configurational degrees of freedom within each structural motif, while preserving the underlying crystal symmetry and topology of the prototype structure.
(ii) Lattice-volume scaling
Lattice-volume scaling is applied to approximate the equilibrium bond lengths for the newly generated compounds. (Fig. 2b). Since the atomic sizes of the elements in the template structure may differ significantly from those in the target system, the equilibrium volumes of the substituted compounds will also differ substantially. To account for this, scaling factors are applied: typically multiple values ranging from 0.90 to 1.10, which uniformly expand or contract the unit cell volume. For example, a prototype with lattice constant $a_{0}$ is scaled to $0.92a_{0}$, $0.96a_{0}$, $1.0a_{0}$, $1.04a_{0}$, and $1.08a_{0}$, as described in [JOSS, FeCoZr, CeCoCu].
@@ 2.2 Rapid stability screening using ML models
To efficiently predict the formation energies and eliminate high-energy candidates, exa-AMD employs an ML model, Crystal Graph Convolutional Neural Network (CGCNN), for high-throughput stability screening [CGCNN]. CGCNN represents each hypothetical crystal as a graph, encoding atoms as nodes (featuring atomic number and chemical environment) and bonds as edges (including interatomic distances). Each hypothetical structure generated from the previous step is evaluated using the CGCNN model that is trained to predict the formation energy. The initial, universal CGCNN model we utilized in this framework was trained based on 28046 structures from Materials Project, as described in [CGCNN]. The mean absolute error of this model is typically around 0.1–0.2 eV/atom.
Down selection is then carried out to remove duplicates and structures with high energies. First, candidates are sorted by the predicted formation energy, and basic filters (a formation energy cutoff, e.g., $E_{\mathrm{f}}<0$ eV/atom, optional element-fraction limits, and a maximum-atoms limit) are applied. To identify unique crystal structures, we then group the remaining candidates by their reduced composition. Within each group, the structures are processed from lowest to highest formation energy. A structure is kept for subsequent DFT calculation only if it does not match any previously kept structure from that group, as determined by the pymatgen.analysis.structure_matcher.StructureMatcher class with the default difference tolerance [pymatgen]. This results in typically 1,000–4,000 unique candidates for first-principles calculations in the next step [FeCoZr, CeCoCu, JOSS].
Our previous work shows that the use of ML models drastically reduces the computational time from months (using brute force DFT) to minutes for over $10^{6}$ structures [CeCoCu, CeFeX, FeCoC, FeCoP, FeCoSi, FeCoZr, LaCoPb, PNAS, JMCA, Sun2022]. This screening enables the rapid exclusion of high-energy or chemically implausible candidates before any quantum calculation is performed.
After the completion of the first round of the framework (i.e., including DFT calculations and post-processing as discussed in Sections 2.3 and 2.4), a system-specific, second-generation CGCNN model is retrained using several hundred to thousands of new DFT-relaxed ground states, reducing the mean absolute errors to as low as 0.03–0.05 eV/atom for the chemical systems of interest [PNAS, JMCA, FeCoZr, CeCoCu, FeCoC]. Key hyperparameters for training include 3–6 convolutional layers, batch size of 64–256, 100–200 epochs, stochastic gradient decent optimization, and 80/10/10 data splits for training/validation/testing, respectively. In this work, we choose to use the default setting of these parameters provided by Xie, et al [CGCNN].
Since the evaluation of each candidate structure is independent, this approach provides a scalable and accurate procedure to select the most promising candidate materials for the costly quantum calculations in the next step. Importantly, our framework allows users to replace or augment the ML model with other ML architectures or retrained models specific to their system of interest, enabling improved prediction accuracy and tailored screening strategies. Such flexibility allows adaptation to a wide range of material classes and accelerates convergence to plausible candidate structures.
@@ 2.3 First-Principles Calculations
Formation energies per atom are always referenced to the relaxed energies of the elemental (and, if needed, binary) phases, and convex hull constructions are used to assess thermodynamic stability—a compound is designated as stable, metastable ($E_{\rm hull}<0.1$ eV/atom), or unstable in this step. For correlated systems, a Hubbard $U$ correction is applied within standard DFT for more accurate valence treatments when necessary. Dynamical stability checks, such as phonon calculations and ab initio molecular dynamics, can be conducted for representative new phases as post-processing steps. By parallelizing DFT jobs across CPU/GPU clusters using Parsl, thousands of structure relaxations with varying size and complexity can be completed within hours to days—enabling practical exploration of otherwise intractable structural and compositional spaces [JOSS, FeCoZr, CeCoCu, PNAS, JMCA]. Although this first-principles calculation stage primarily uses VASP as the default DFT engine, the workflow is designed to be agnostic about the choice of DFT software. Depending on user preferences and available computational resources, alternative software packages such as Quantum ESPRESSO [QEa, QEb] can be used. This modularity not only offers flexibility in computation but also enables benchmarking and method comparison to ensure robust validation of candidate materials.
@@ 2.4 Convex Hull and Thermodynamic Stability
$E_{\text{form}}=E_{\text{tot}}(\text{compound})-\sum_{i}n_{i}E_{\text{ref}}(i),$ (1)
@@
Candidates are compared against all known phases to compute their energy above the convex hull ($E_{\mathrm{hull}}$). To construct the reference convex hull, we retrieve all thermodynamically stable and metastable structures for the relevant chemical system (e.g., Fe-Co-Zr ternary system) from the Materials Project database [MP] using their Python Application Programming Interface (API). Specifically, we use the MPRester.get_entries_in_chemsys() method to obtain all ComputedStructureEntry objects containing the target elements, which include elemental phases (e.g., pure Fe, Co, Zr), binary compounds (e.g., Fe-Co, Fe-Zr, Co-Zr), and known ternary phases in the chemical system. Each entry contains the DFT-calculated total energy and crystal structure, from which formation energies are computed relative to the elemental reference states. These formation energies, combined with our newly calculated DFT results for the generated candidate structures, are used to construct the convex hull — the set of thermodynamically stable phases, where any compound above this hull is metastable or unstable [convexhull]. Compounds on the hull ($E_{\mathrm{hull}}=0$ eV/atom) are identified as thermodynamically stable, while those near the hull (e.g., $E_{\mathrm{hull}}<0.1$ eV/atom) are considered metastable and potentially synthesizable under non-equilibrium conditions. This rigorous convex hull analysis is essential for predicting which compounds can potentially be synthesized in practice, as confirmed by literature and data-mined studies [AFLOW, GNOME]. After all candidate structures complete DFT relaxation and their formation energies are computed, the convex hull is updated by combining the newly calculated structures with the retrieved existing known phases. Newly discovered low-energy structures may shift the convex hull downward, potentially reclassifying previously stable Materials Project phases as metastable, while simultaneously establishing new stable or near-stable phases. The final output represents the most complete and up-to-date phase diagram for the chemical system, incorporating both the Materials Project reference data and all validated compounds from the current discovery campaign.
It is important to note that the scope of the predicted phases and phase diagrams generated by exa-AMD is fundamentally determined by the set of structure prototypes supplied at the beginning of the workflow. These initial prototypes, either provided by default with exa-AMD or customized by the user, define the structural motifs accessible to elemental substitution, combinatorial screening, and all subsequent ML and DFT evaluations. Consequently, all new compounds, low-energy structures, and phase diagram updates produced by the current version of exa-AMD are restricted to the chemical and structural space spanned by these templates. While this approach enables high throughput calculations and chemical flexibility, it does not guarantee exhaustive exploration of all possible structure types. As demonstrated in our previous work on the Fe-Co-C system [FeCoC], this limitation can be addressed by integrating adaptive genetic algorithms (AGA) [AGA1, AGA2] capable of exploring and predicting new structure types that are absent from existing databases or user-supplied templates. In future releases of exa-AMD, we plan to natively incorporate AGA-based structure prediction to overcome the prototype bottleneck and expand the framework’s reach to uncover genuinely novel structural motifs and compounds beyond the current prototype-driven paradigm.
@@ 3 Software Implementation
exa-AMD is implemented in Python and designed for clarity, portability, and extensibility. It encompasses plug-in scripts for structure handling, database access, machine learning inference, and DFT job submission. The use of the Parsl library [Parsl] ensures efficient, scalable workflow orchestration involving task parallelism, fault tolerance, and elastic resource management. Users can configure resource allocation, execution platforms, and job scheduling via flexible configuration files, allowing seamless portability from local workstations to large-scale HPC systems. Key features of the software implementation include:
@@
    Modular design: The framework’s modular architecture allows users to customize or replace key components to adapt to diverse research needs. For example, users can provide their own customized initial structure pool as discussed above. Moreover, the machine learning model (default is CGCNN currently), can be replaced by other advanced models with the same interface, such as ALIGNN [ALIGNN], M3GNet[M3GNET], MEGNet[MEGNET], CHGNet[CHGNET], etc. The first-principles component also supports multiple DFT engines such as VASP, Quantum ESPRESSO, and others.
@@
    I/O and database integration: Structures are sourced from our exa-AMD dataset [exaAMD-dataset], curated from multiple databases or user-provided datasets, with outputs systematically organized into logs, results, and configuration files to ensure comprehensive data provenance and reproducibility.
@@
    Parallel execution: High-throughput tasks like ML predictions and DFT calculations are automatically parallelized, with configurable resource allocations (node types and counts) for elastic scaling on supercomputers and clouds.
@@
At the heart of exa-AMD’s scalability and automation is the use of Parsl [Parsl], a flexible, Python-based parallel programming and workflow library designed to efficiently orchestrate scientific pipelines across heterogeneous computational resources. Parsl enables exa-AMD to decompose each major workflow step—structure generation, ML-based screening, structure similarity filtering, DFT calculations, and property post-processing—into fine-grained, asynchronous tasks (“apps” in Parsl terminology) that can be dynamically mapped onto available compute resources such as CPUs and GPUs.
Building on this task-based execution model, users provide a Parsl configuration that operates independently of the scientific workflow itself. The configuration defines how each workflow stage is mapped onto computational resources. Fig. 3 shows an example mapping on a heterogeneous machine where CPU nodes are dual-socket and GPU nodes contain four accelerators each. The figure illustrates several key aspects of exa-AMD’s execution model.
Figure 3: Example mapping of exa-AMD stages to hardware via Parsl executors: CPU nodes (blue) and GPU nodes (green).
First, at every stage of the workflow, exa-AMD achieves scalability through task-level concurrency across nodes and intra-node parallelism. For example, independent tasks, such as VASP calculations on different candidate structures, are executed concurrently across multiple nodes, while multithreading and GPU acceleration exploit parallelism within each node.
Second, certain stages (structure generation and formation-energy prediction) are directly parallelized through Parsl. Users can specify how the overall workload is partitioned into independent chunks, which determines the number of concurrent tasks. In the example shown in Fig. 3, the workload is divided into four chunks, producing four parallel tasks.
Third, the execution model does not require dedicating one node per task. Multiple lightweight tasks can share the computational resources of a single node. For instance, four independent tasks may run concurrently on a single GPU node. The user determines the allocation strategy according to the size and complexity of the target compound, balancing throughput against resource efficiency. In practice, the DFT stage dominates the overall computational cost (see Section 5), which is reflected in the figure by allocating sixteen GPU nodes to this stage.
Fourth, the post-processing stage can reuse the existing GPU allocation to avoid additional wait time in the queue on a cluster. In this stage, exa-AMD performs VASP single-point total-energy calculations on GPUs, followed by short CPU-only tasks such as computing and plotting the convex hull. For such brief operations, reusing the same GPU allocation is generally more efficient than requesting a separate CPU allocation, as it eliminates unnecessary scheduling delays.
In addition, for performance reasons, the number of concurrent tasks being processed at a time, used in the ML prediction stage should be an integer divisor of that used in the structure generation stage, because the initial partitioning of the problem is established at the first stage. In the example of Fig. 3, the initial workload is divided into four chunks and thus produces four tasks. Structure generation runs on four nodes with full concurrency, while the ML prediction stage, using only two nodes, executes the same four tasks in two successive waves of two concurrent jobs.
@@
Parsl also enables robust resumability and error handling within exa-AMD. Every task’s state and results are tracked, allowing failed calculations to be retried or workflows to resume from intermediate stages in the event of an interruption. This highly modular and fault-tolerant approach not only maximizes resource usage and throughput on shared HPC queues but also supports rapid development and reproducibility—key requirements for cutting-edge, community-driven materials discovery frameworks.
@@
By decoupling workflow logic from execution configuration, exa-AMD with Parsl facilitates portability. Users can port identical workflows across computing systems by simply adjusting the configuration files, making rapid prototyping, exploration, and full production campaigns equally straightforward.